THE SUBMILLIMETER SPECTRUM OF GLYCOLALDEHYDE FROM 500 GHz TO 1.2 THz

\underline{P. BRANDON CARROLL}, Department of Chemistry, Emory University, Atlanta, GA 30322; BRIAN J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA, 91109; and SUSANNA L. WIDICUS WEAVER, Department of Chemistry, Emory University, Atlanta, GA 30322.

Glycolaldehyde is the simplest α-hydroxy aldehyde and the only sugar-related molecule definitively detected in the interstellar medium to date. Previous spectral studies^{a}^{,}^{b} have reported its rotational spectrum up to ∼350 GHz. With the advent of several new submillimeter and THz astronomy facilities equipped with high-resolution spectrometers, it is important to begin collecting and analyzing higher frequency spectra for all known interstellar molecules. We have acquired the submillimeter/THz direct absorption spectrum of glycolaldehyde from 520 – 600 GHz and from 1.1 – 1.2 THz. The spectra were measured at JPL using a 3-meter static cell in double-pass configuration. The submillimeter radiation was generated using a standard microwave frequency synthesizer coupled with cascaded frequency multipliers. A Schottky diode detector was used during acquisition of the lower frequency spectra, and an Si hot electron bolometer detector was used for scans above 1 THz. The analysis of these spectra is underway, and new line assignments are being combined with the assignments from previous studies to determine a global spectral fit. The details of the experiment and progress on the spectral analysis will be reported.

aMarstokk, K. M.; Møllendal, H., J. Mol. Struct., 5, 205-213 (1970).

bWidicus Weaver, S. L. et al., Astrophys. J. Suppl., 158, 188-192 (2005).